In atomic structure how can be bohr be proved right in accordance with heisenberg uncertainity principle?

Grade:10

2 Answers

Godfrey Classic Prince
12 years ago

Dear Ramendra Ck,

Heisenberg's Uncertainity principle says that it is impossible to measure simultaneously both the postion & velocity(or momentum) of a sub-microscopic particle with absolute accuracy or certainity .

But any one of the two , that is either position or velocity of a sub-microscopic particle(electron) can be predicted accurately at a given time !

Thus Bohr derived the formula for the velocity of an Electron in an nth orbital to be

By Bohr's equation: L = nh /(2 pi)
Plugging into the general formula:
nh/(2 pi) = mvr
v = nh / ( 2pi)(mr)
Substituting the values the velocity of the electron in the nth orbital is given to be
v = 2.19*106 Z/n 
Z- Atomic Number
n- Orbital Number
Hope thsi helped you immensely ..!!
If you loved the answer then please do not forget to approve it by clicking on "Yes" below..!!!
Chetan Mandayam Nayakar
12 years ago

Strictly speaking Bohr's theory is wrong. It belongs to what is called the "old quantum theory". One must use Schrodinger's equation and either wave mechanics or Dirac notation. It is well beyond the scope of IITJEE.